Grade 10Algebra

A pair of dice is rolled until a total of 5 or7 is obtained. Find the probabilty that a total of 5 comes before a total of 7.

harshal  devikar
13 Years agoGrade 10

1 Answer

Vikas TU
13 Years ago

p(total 5) = 4/36
p(total 7) = 6/36
p(total 5 and total 7) = 0
p(total 5 or total 7) = p(total 5) + p(total 7) - p(total 5 and total 7)
=4/36 + 6/36 = 10/36 = 5/18

 

plz approve!